# How Far Are We From AGI?

### v1

@@ 4.3 Large-scale Training
Many important techniques have been widely adopted in popular DL frameworks to fit larger models into devices with fixed memory. CPU offloading allows models to selectively transfer weights (layers) or KV cache to CPU with more memory, and only load essential network parts to GPU for processing. When pushed to the extreme, FlexGen (Sheng et al., 2023b) can achieve significant batch throughput of OPT-175B on a single 16GB GPU. Gradient checkpointing (Chen et al., 2016) reduces peak memory usage by recomputing parts of the computational graph during back-propagation. There is no doubt that efficient memory management will remain as the core investment direction that enables the deployment of scalable systems and parallel processing of larger batches.
@@
Efficient Fine-tuning Pretrained large models often internalize a tremendous amount of knowledge which can be unleashed by (instruct) fine-tuning. However, despite the fact that often a relatively small amount of examples are sufficient for successful fine-tuning, the cost and time for doing so are prohibitive and not economical. The main objective for efficient fine-tuning is to figure out a balance between the cost (implementation difficulty, data requirement, training budget, etc) and the performance gap from continual pretraining. A series of parameter-efficient fine-tuning (PEFT) techniques have been developed to meet this challenge, which only requires training a small number of new parameters and often achieves better performance than in-context learning. LoRA (Sheng et al., 2023a) as one of the most popular PEFT methods, draws great attention these days. LoRA and many of its variants (LoHA (Hyeon-Woo et al., 2023), AdaLoRA (Zhang et al., 2023a), Q-LoRA (Dettmers et al., 2023), and recent PiSSA Meng et al. (2024)) insert learnable matrices that are low-rank decompositions of the delta weight matrices. LLaMA-Adapter (Zhang et al., 2023d) employs a learnable adaptation prompt prepended to the input prompt and uses a zero-initialized attention mechanism with zero gating for efficient fine-tuning. Sharing a similar flavor to LoRA, $\text{IA}^{3}$ (Liu et al., 2022b) scales the model activations by learned vectors instead of matrices. Other PEFT methods that insert learnable components showcase strong generalization ability, and prompt-based methods like soft prompting (Lester et al., 2021) add extra learnable parameters to the input embeddings while keeping the original model weights frozen. Adapters (Houlsby et al., 2019) add trainable parameters inside the attention blocks, while Prefix tuning (Li and Liang, 2021) appends learnable vectors to the KV representations in attention. Unlike the traditional PEFT techniques, Zhao et al. (2023c); Basu et al. (2023) first discovered tuning the Layernorm layer of transformers yields decent performance in these models as an unexpectedly strong baseline. Other than twitching model parameters in the PEFT process, GaLore (Zhao et al., 2024b) proposes to apply the LoRA tuning paradigm on the model gradients, and REFT (Wu et al., 2024a) chooses to place a linearly probing strategy between the source model parameters and the optimization goal.
@@
Decentralization There are many works that focus on utilizing dis-aggregated and hardware heterogeneous computing devices over the cloud for model training and inference. One challenge of geographically separated clusters is the communication overhead, which makes data movement costly (training data, gradients, KV cache, etc.) and eclipses decentralization’s benefits. CacheGen (Liu et al., 2023c) compresses the KV cache with an encoder into compact bitstream representations, reducing the latency for context fetching and processing. CocktailSGD (Wang et al., 2023g) employs a combination of sparsification and quantization techniques, which makes fine-tuning LLMs up to 20B size with slow networks possible with only minimal slowdown compared to data center’s fast interconnect. Collaborative training crowdsources commodity GPUs from individual users, the most prominent example of which is Petals (Borzunov et al., 2022), a system capable of serving and fine-tuning BLOOM-176B and OPT-175B with decent performance (e.g. supports interactive sessions) using only mediocre GPUs from multiple parties. Decentralized AI systems open up the possibility of bridging devices across the globe, which ensures fault-tolerance (Ryabinin et al., 2023) and compatibility of heterogeneous devices plus networks (Jiang et al., 2023e; Yuan et al., 2023), as well as optimizing limited network bandwidth (Wang et al., 2023i) and data privacy (Tang et al., 2023).
@@
Training Dynamics & Scaling The science of large language models is mysteriously difficult to grasp, the understanding of which can drastically improve the development of various AIs. However, most successful LLMs are not fully “open” not just in terms of data and model weights but other aspects such as the intermediate checkpoints and artifact logging that can assist in reasoning about the training dynamics as we scale models to different sizes. (Xia et al., 2023a) analyzes the intermediate checkpoints of OPT models (Zhang et al., 2022a) on various downstream tasks, which attempts to emphasize the perplexity as a predictive indicator of a model’s performance than its size, showing that larger models hallucinate less often and that models tend to exhibit minimal return during early stage of the training. Complimentary to this, (Tirumala et al., 2022) focuses on studying different memorization capabilities across the model size, dataset size, and learning rate and proposes an interesting hypothesis on the importance of nouns and numbers as the unique identifier for memorizing individual training examples. Besides pure analysis, Pythia (Biderman et al., 2023) introduces a suite of 16 LLMs trained on public data, ranging in size from 70M to 12B parameters. With these intermediate checkpoints released to the broader community, it becomes way easier and more efficient for researchers to find answers to questions related to training dynamics by examining and benchmarking individual saved weights and losses. Finally, OLMo (Groeneveld et al., 2024), on top of that, graciously release the whole framework, including training data and training and evaluation code, for the benefit of making the study of the science behind LLMs easier.
@@ 4.4 Inference Techniques
AGI inference systems need to ensure user responsiveness, availability, and efficiency, which helps unleash the ultimate potential of large models from the training phase and revolutionize how users interact with the system. Hence, in this section, we give an overview of several techniques that try to accelerate auto-regressive decoding, balance request scheduling, and serve a massive number of models with different capabilities in the cluster, which will inspire future system effort across the spectrum.
Decoding Algorithm In this paper, we focus mostly on exact decoding acceleration where we want to maximize the performance while staying faithful to the original model without compromising the accuracy. (Miao et al., 2023) gives a comprehensive review of several approximate methods such as sampling strategies, non-autoregressive decoding, semi-autoregressive decoding, block-parallel decoding, and early existing, etc. A large body of works explore the idea of speculative decoding (Leviathan et al., 2023) with the central idea of trading parallel computation for higher chances of generating multiple tokens at once. Usually, a speculative decoding process starts with an efficient draft model that makes predictions of multiple steps, the resulting proposals verified by the target model we want to sample. However, there are many challenges involved, including 1) how to make the draft model lightweight while still generating useful guesses for efficient progress, 2) how to avoid extensive architecture change and fine-tuning for faster adaptation, and 3) how to deploy the draft model more effectively. The simplest yet effective variant is called Prompt-lookup Decoding (Saxena, 2023) where the draft model is replaced by simple prefix string matching from an existing database for generating candidate tokens. This model-agnostic approach can decode extremely fast without any fine-tuning or model change, but the performance heavily depends on the quality and diversity of the string pool. To facilitate faster verification over a large number of candidates, SpecInfer (Miao et al., 2024) organizes the outputs of the draft models into a token tree, with each node being a candidate token, the correctness of which can be efficiently checked in parallel by the base model. Following a similar idea, Medusa (Cai et al., 2024a) introduces a tree attention mechanism to simultaneously check all tokens from the medusa heads, which is realized by a special mask pattern for efficient parallel computation. Self-speculative decoding (Zhang et al., 2023k) proposes to completely discard the requirement of a draft model and generate candidate sequences by selectively skipping a subset of intermediate layers.
@@
Request Scheduling Request scheduling for LLMs poses several unique challenges compared to traditional machine learning systems with structured inputs. Some important features for a mature request scheduling strategy include 1) efficient pre-fetching of the context (user information, past KV cache, and model adapter, etc) for a given input, 2) handling examples with variable sequence lengths for maximal GPU utilization, and 3) trading-off various request-level metrics such as time-to-first-token (TTFT), job completion time (JCT), batch token throughput, and inference latency. Orca (Yu et al., 2022) proposed an iteration-level scheduling mechanism to meet the auto-regressive nature of LLM inference requests, which, when coupled with a technique called selective batching for better hardware utilization, outperforms previous inference engines like FastTransformer (NVIDIA, 2023a) in terms of throughput and latency. Other strategies of dynamic batching are explored extensively, such as the continuous batching from vLLM (Kwon et al., 2023a) and in-flight batching from TensorRT-LLM (NVIDIA, 2023b) are explored. Rather than request-level scheduling, FastServe (Wu et al., 2023c) exploits the autoregressive pattern of LLM inference to enable preemption at the granularity of each output token, which optimizes JCT with a novel skip-join Multi-Level Feedback Queue scheduler that leverages the information of input lengths for better efficiency. The inference workload is strongly tied to the average sequence lengths of examples, and hence, we want to minimize the gap between the longest and shortest sentences. $S^{3}$ (Jin et al., 2023a) predicts the potential response length for each example in the batch, which is used for fitting more examples under the same memory constraint (e.g. GPU memory). Dynamic SplitFuse from DeepSpeed-FastGen (Holmes et al., 2024) takes the insight of LLM inference (the consequence of changing batch size v.s. number of tokens on model’s performance) and proposes a token composition strategy. Dynamic SplitFuse runs at a consistent forward size by taking partial tokens from prompts and composing this with generation. For example, long prompts are split into smaller chunks across several forward iterations, and short ones are composed to align with the other requests. With this strategy, the system not only provides better efficiency and responsiveness but also reduces the variance over requests.
@@
Multi-model Serving Besides serving multiple replicas of the same model, being capable of deploying numerous task-specialized models efficiently becomes an important feature for many application scenarios (LLM agents, persona chat-bots, privacy-sensitive assistants, etc.). Naively scaling the number of instances, however, is both computationally prohibitive and resource-wasteful. With the advancement of PEFT techniques, serving a base model with diversified adapters becomes a paradigm favored by many practitioners due to the fact that PEFT models are light-weight and easy to maintain while being flexible and powerful. The major challenge for multi-model (PEFT) serving is how to dynamically and efficiently load the “right” ( measured by latency or task performance, etc) adapter for each example. Punica (Chen et al., 2023h) enables the efficient computation of heterogeneous LoRA heads in a batch with a newly designed CUDA kernel that shares a single pre-trained model, achieving up to 12×12$\times$12 × higher throughput while only adding slight extra latency. S-LoRA (Sheng et al., 2023a) introduces Unified Paging which uses a unified memory pool for dynamic adapter management and highly optimized CUDA kernels for parallelizing LoRA computation. LoRAX (Predibase, 2023) additionally provides adapter exchange scheduling which asynchronously prefetches and offloads adapters between GPU and CPU memory and schedules request batching to optimize the aggregated throughput. With these systems, it becomes possible to serve over a thousand different LoRA heads on a single GPU, opening up a broader possibility such as model collaboration, task-generalization, and model merging.
@@ 4.5 Cost and Efficiency
Data Economy Data plays a pivotal role in a model’s performance and the question of how much data value is fundamentally important for many reasons: 1) what data should we collect to add to the existing data mix for improving performance 2) how should we reasonably pay for data provider, and 3) can we remove non-essential data (outliers) to make our models more robust. To answer these questions, many works from computer science and economics (game theory) have explored different formalisms to define what “data value” means and how to estimate it efficiently. Shapley value comes in handy from the classic game theory, which uniquely satisfies several natural properties of equitable data valuation (Ghorbani and Zou, 2019). Naive computation of Data Shapley requires exponential time, and hence Monte Carlo (Ghorbani and Zou, 2019) and gradient-based methods are used to make it efficient (Jia et al., 2019). TracIn takes a similar idea of tracing the influence of individual training examples with gradient information. To make these algorithms practical and easy to use, DataScope (Karlaš et al., 2022) is developed as an end-to-end system that can efficiently compute the Shapley value of training data over the whole pipeline consisting of various ML algorithms and data transformation, making it a powerful tool for data debugging. With more mature data valuation, data providers are more motivated to contribute, fostering a more healthy and robust data-centric ecosystem.
@@
Model Combination Model combination (MC) strives to improve the overall system’s performance by either orchestrating or merging a series of (specialized) large models. The key benefits of model combination rely on the fact that there is usually little or no need for explicit training, and they can often result in better down-stream performance and task-generalization capability. FrugalGPT (Chen et al., 2023i) routes quests in a cascading manner to different LLMs and uses a learned scoring function to decide whether to return the intermediate results in a flexible way, which drastically lowers the cost and improves the quality. Merging weights of multiple LLMs have been explored in many forms and shown to be effective. Popular methods include simple averaging (Wortsman et al., 2022), task arithmetic (Ilharco et al., 2023), multi-modal (encoders) merging (Wu et al., 2023b; Sung et al., 2023), merging based on learned routing function (Lu et al., 2023), SLERP, and weighted (conjugate gradient descent (Tam et al., 2023), stochastic and population-based optimization algorithm (Huang et al., 2024)) merging. MC can also be extremely promising for federated learning because only model weights are exchanged, and hence, data privacy is easier to guarantee. CoID Fusion (Don-Yehiya et al., 2023) proposes to collaboratively improve the multi-task learning of a base model by sending copies to workers and fusing the learned weights without data communication. The model combination can lead to compound systems [^4] https://bair.berkeley.edu/blog/2024/02/18/compound-ai-systems which consists of multiple LLMs working in synergy (via merging, routing, or knowledge sharing). AIOS (Mei et al., 2024) devises a mechanism to integrate multiple LLM agents into an operating system, the synergistic combination of which enables increasingly complex, multi-modal tasks that require reasoning, execution, and interaction with the physical world. Tandom Transformer (S et al., 2024) equips a smaller less accurate model with attention to the rich representation of a larger model that can process multiple tokens simultaneously, which serves as a stitched student-teacher system that improves both accuracy and efficiency in the downstream tasks. Developing complex compound systems poses several challenges: 1) how to co-optimize multiple LLMs, 2) identifying the failing (insecure) component is a lot harder than debugging a monolithic system, and 3) how to design mature data pipelines for different components of the large system. Addressing (some of) these problems will significantly increase the possibility of exciting new applications.
@@
Automation The increasing complexity of building a large foundation model requires a more mature automation process for democratization and agile development. AutoML (Hutter et al., 2019) has achieved remarkable success in many machine learning tasks over the last couple of years (Zimmer et al., 2020; Feurer et al., 2020; Erickson et al., 2020), which proves itself as a promising solution for large model automation. Applying AutoML techniques to LLMs, however, poses many challenges such as the cost for pretraining, the multitude of different stages, and performance indicators, making holistic optimization difficult or even infeasible (Tornede et al., 2024). Nonetheless, we will introduce some exemplary attempts targeting certain stages of the whole system. PriorBand (Mallik et al., 2023) tries to bridge the gap in the cost of Hyperparameter Optimization (HPO) between traditional ML and modern DL via utilizing expert beliefs and cheap proxy tasks. AdaBERT (Chen et al., 2021a) is an automated task-specific compression algorithm based on differentiable Neural Architecture Search (NAS) which is guided by both a task-oriented knowledge distillation loss and an efficiency-aware loss. To reduce the burden of prompt engineering, Automatic Prompt Engineer (APE) (Zhou et al., 2023c) proposes to leverage the interplay of several LLMs for automatic prompt generation and selection where one LLM proposes or modifies a prompt and another LLM rates it for selection. EcoOptiGen (Wang et al., 2023f) optimizes the utility and cost of decoding by finding better hyper-parameters, such as the number of responses, temperature, and max tokens, which demonstrates the potential of applying AutoML for the inference stage. One extremely exciting approach for building complicated and compound systems is to ask multiple LLMs to solve a big problem in a decomposed way cooperatively. One realization is to prompt the LLM or VLM to serve different purposes in a pipeline, which can be tremendously challenging to tune, optimize, modularize, and debug. DSPy (Khattab et al., 2023; 2022) tackles this by first separating the flow of the system from the parameters (i.e., model prompts and weights) at each step, and then dedicated algorithms are used to tune them with user’s defined metric. Even with all the works discussed above, the integration and development of Automated Large Models (AutoLM) has many challenges and opportunities simultaneously.

### v2

@@ 4.3 Large-scale Training
Many important techniques have been widely adopted in popular DL frameworks to fit larger models into devices with fixed memory. CPU offloading allows models to selectively transfer weights (layers) or KV cache to CPU with more memory, and only load essential network parts to GPU for processing. When pushed to the extreme, FlexGen (Sheng et al., 2023b) can achieve significant batch throughput of OPT-175B on a single 16GB GPU. Gradient checkpointing (Chen et al., 2016) reduces peak memory usage by recomputing parts of the computational graph during back-propagation. There is no doubt that efficient memory management will be remained as the core investment direction that enables the deployment of scalable systems and parallel processing of larger batches.
@@
Efficient Fine-tuning Pretrained large models often internalize a tremendous amount of knowledge which can be unleashed by (instruct) fine-tuning. However, despite the fact that often a relatively small amount of examples are sufficient for successful fine-tuning, the cost and time for doing so are prohibitive and not economical. The main objective for efficient fine-tuning is to figure out a balance between the cost (implementation difficulty, data requirement, training budget, etc) and the performance gap from continual pretraining. A series of parameter-efficient fine-tuning (PEFT) techniques have been developed to meet this challenge, which only requires training a small number of new parameters and often achieves better performance than in-context learning. LoRA (Sheng et al., 2023a) as one of the most popular PEFT methods, draws great attention these days. LoRA and many of its variants (LoHA (Hyeon-Woo et al., 2023), AdaLoRA (Zhang et al., 2023a), Q-LoRA (Dettmers et al., 2023), and recent PiSSA Meng et al. (2024)) insert learnable matrices that are low-rank decompositions of the delta weight matrices. LLaMA-Adapter (Zhang et al., 2023d) efficiently fine-tunes LLaMA into an instruction-following model with very little computational budget. A set of learnable adaptation prompts are first prepended to the context and they train a zero-initialized attention mechanism with zero gating with only 52K self-instruct demonstrations. The resulting extra 1.2M parameters from the adapter can give high-quality outputs, comparable to fully fine-tuned results. Sharing a similar flavor to LoRA, $\text{IA}^{3}$ (Liu et al., 2022b) scales the model activations by learned vectors instead of matrices. Other PEFT methods that insert learnable components showcase strong generalization ability, and prompt-based methods like soft prompting (Lester et al., 2021) add extra learnable parameters to the input embeddings while keeping the original model weights frozen. Adapters (Houlsby et al., 2019) add trainable parameters inside the attention blocks, while Prefix tuning (Li and Liang, 2021) appends learnable vectors to the KV representations in attention. Unlike the traditional PEFT techniques, Zhao et al. (2023c); Basu et al. (2023) first discovered tuning the Layernorm layer of transformers yields decent performance in these models as an unexpectedly strong baseline. Other than twitching model parameters in the PEFT process, GaLore (Zhao et al., 2024b) proposes to apply the LoRA tuning paradigm on the model gradients, and REFT (Wu et al., 2024a) chooses to place a linear probing strategy between the source model parameters and the optimization goal.
@@
Decentralization Many works focus on utilizing dis-aggregated and hardware heterogeneous computing devices over the cloud for model training and inference. One challenge of geographically separated clusters is the communication overhead, which makes data movement costly (training data, gradients, KV cache, etc.) and eclipses decentralization’s benefits. CacheGen (Liu et al., 2023c) compresses the KV cache with an encoder into compact bitstream representations, reducing the latency for context fetching and processing. CocktailSGD (Wang et al., 2023g) employs a combination of sparsification and quantization techniques, which makes fine-tuning LLMs up to 20B size with slow networks possible with only minimal slowdown compared to data center’s fast interconnect. DiLoCo (Douillard et al., 2023) introduces a novel federated averaging algorithm run on islands of devices that are poorly connected, which claims to perform as well as fully synchronous optimization on C4 datasets while communicating 500 times less. Collaborative training crowdsources commodity GPUs from individual users, the most prominent example of which is Petals (Borzunov et al., 2022), a system capable of serving and fine-tuning BLOOM-176B and OPT-175B with decent performance (e.g. supports interactive sessions) using only mediocre GPUs from multiple parties. Decentralized AI systems open up the possibility of bridging devices across the globe, which ensures fault-tolerance (Ryabinin et al., 2023) and compatibility of heterogeneous devices plus networks (Jiang et al., 2023e; Yuan et al., 2023), as well as optimizing limited network bandwidth (Wang et al., 2023i) and data privacy (Tang et al., 2023).
@@
Training Dynamics & Scaling The science of large language models is mysteriously difficult to grasp, the understanding of which can drastically improve the development of various AIs. However, most successful LLMs are not fully “open” not just in terms of data and model weights but other aspects such as the intermediate checkpoints and artifact logging that can assist in reasoning about the training dynamics as we scale models to different sizes. (Xia et al., 2023a) analyzes the intermediate checkpoints of OPT models (Zhang et al., 2022a) on various downstream tasks, which attempts to emphasize the perplexity as a predictive indicator of a model’s performance than its size, showing that larger models hallucinate less often and that models tend to exhibit minimal return during early stage of the training. Complimentary to this, (Tirumala et al., 2022) focuses on studying different memorization capabilities across the model size, dataset size, and learning rate and proposes an interesting hypothesis on the importance of nouns and numbers as the unique identifier for memorizing individual training examples. Besides pure analysis, Pythia (Biderman et al., 2023) introduces a suite of 16 LLMs trained on public data, ranging in size from 70M to 12B parameters. With these intermediate checkpoints released to the broader community, it becomes way easier and more efficient for researchers to find answers to questions related to training dynamics by examining and benchmarking individual saved weights and losses. Finally, OLMo (Groeneveld et al., 2024), on top of that, graciously releases the whole framework, including training data and training and evaluation code, for the benefit of making the study of the science behind LLMs easier.
@@ 4.4 Inference Techniques
AGI inference systems need to ensure user responsiveness, availability, and efficiency, which helps unleash the ultimate potential of large models from the training phase and revolutionize how users interact with the system. Hence, in this section, we give an overview of several techniques that try to accelerate auto-regressive decoding, balance request scheduling, and serve a massive number of models with different capabilities in the cluster, which will inspire future system efforts across the spectrum.
Decoding Algorithm In this paper, we focus mostly on exact decoding acceleration where we want to maximize the performance while staying faithful to the original model without compromising the accuracy. (Miao et al., 2023) gives a comprehensive review of several approximate methods such as sampling strategies, non-autoregressive decoding, semi-autoregressive decoding, block-parallel decoding, and early existing, etc. A large body of works explores the idea of speculative decoding (Leviathan et al., 2023) with the central idea of trading parallel computation for higher chances of generating multiple tokens at once. Usually, a speculative decoding process starts with an efficient draft model that makes predictions of multiple steps, the resulting proposals verified by the target model we want to sample. However, there are many challenges involved, including 1) how to make the draft model lightweight while still generating useful guesses for efficient progress, 2) how to avoid extensive architecture change and fine-tuning for faster adaptation, and 3) how to deploy the draft model more effectively. The simplest yet effective variant is called Prompt-lookup Decoding (Saxena, 2023) where the draft model is replaced by simple prefix string matching from an existing database for generating candidate tokens. This model-agnostic approach can decode extremely fast without any fine-tuning or model change, but the performance heavily depends on the quality and diversity of the string pool. To facilitate faster verification over a large number of candidates, SpecInfer (Miao et al., 2024) organizes the outputs of the draft models into a token tree, with each node being a candidate token, the correctness of which can be efficiently checked in parallel by the base model. Following a similar idea, Medusa (Cai et al., 2024a) introduces a tree attention mechanism to simultaneously check all tokens from the medusa heads, which is realized by a special mask pattern for efficient parallel computation. Self-speculative decoding (Zhang et al., 2023k) proposes to completely discard the requirement of a draft model and generate candidate sequences by selectively skipping a subset of intermediate layers.
@@
Request Scheduling Request scheduling for LLMs poses several unique challenges compared to traditional machine learning systems with structured inputs. Some important features for a mature request scheduling strategy include 1) efficient pre-fetching of the context (user information, past KV cache, and model adapter, etc) for a given input, 2) handling examples with variable sequence lengths for maximal GPU utilization, and 3) trading-off various request-level metrics such as time-to-first-token (TTFT), job completion time (JCT), batch token throughput, and inference latency. Orca (Yu et al., 2022) proposed an iteration-level scheduling mechanism to meet the auto-regressive nature of LLM inference requests, which, when coupled with a technique called selective batching for better hardware utilization, outperforms previous inference engines like FastTransformer (NVIDIA, 2023a) in terms of throughput and latency. Other strategies of dynamic batching are explored extensively, such as the continuous batching from vLLM (Kwon et al., 2023a) and in-flight batching from TensorRT-LLM (NVIDIA, 2023b) are explored. Rather than request-level scheduling, FastServe (Wu et al., 2023c) exploits the autoregressive pattern of LLM inference to enable preemption at the granularity of each output token, which optimizes JCT with a novel skip-join Multi-Level Feedback Queue scheduler that leverages the information of input lengths for better efficiency. The inference workload is strongly tied to the average sequence lengths of examples, and hence, we want to minimize the gap between the longest and shortest sentences. $S^{3}$ (Jin et al., 2023b) predicts the potential response length for each example in the batch, which is used for fitting more examples under the same memory constraint (e.g. GPU memory). Dynamic SplitFuse from DeepSpeed-FastGen (Holmes et al., 2024) takes the insight of LLM inference (the consequence of changing batch size v.s. number of tokens on model’s performance) and proposes a token composition strategy. Dynamic SplitFuse runs at a consistent forward size by taking partial tokens from prompts and composing this with generation. For example, long prompts are split into smaller chunks across several forward iterations, and short ones are composed to align with the other requests. With this strategy, the system not only provides better efficiency and responsiveness but also reduces the variance over requests.
@@
Multi-model Serving Besides serving multiple replicas of the same model, being capable of deploying numerous task-specialized models efficiently becomes an important feature for many application scenarios (LLM agents, persona chat-bots, privacy-sensitive assistants, etc.). Naively scaling the number of instances, however, is both computationally prohibitive and resource-wasteful. With the advancement of PEFT techniques, serving a base model with diversified adapters becomes a paradigm favored by many practitioners due to the fact that PEFT models are lightweight and easy to maintain while being flexible and powerful. The major challenge for multi-model (PEFT) serving is how to dynamically and efficiently load the “right” ( measured by latency or task performance, etc) adapter for each example. Punica (Chen et al., 2023h) enables the efficient computation of heterogeneous LoRA heads in a batch with a newly designed CUDA kernel that shares a single pre-trained model, achieving up to 12×12$\times$12 × higher throughput while only adding slight extra latency. S-LoRA (Sheng et al., 2023a) introduces Unified Paging which uses a unified memory pool for dynamic adapter management and highly optimized CUDA kernels for parallelizing LoRA computation. LoRAX (Predibase, 2023) additionally provides adapter exchange scheduling which asynchronously prefetches and offloads adapters between GPU and CPU memory and schedules request batching to optimize the aggregated throughput. With these systems, it becomes possible to serve over a thousand different LoRA heads on a single GPU, opening up a broader possibility such as model collaboration, task-generalization, and model merging.
@@ 4.5 Cost and Efficiency
Data Economy Data plays a pivotal role in a model’s performance and the question of how much data value is fundamentally important for many reasons: 1) what data should we collect to add to the existing data mix for improving performance 2) how should we reasonably pay for data provider, and 3) can we remove non-essential data (outliers) to make our models more robust. To answer these questions, many works from computer science and economics (game theory) have explored different formalisms to define what “data value” means and how to estimate it efficiently. Shapley value comes in handy from the classic game theory, which uniquely satisfies several natural properties of equitable data valuation (Ghorbani and Zou, 2019). Due to its rich theoretical results, Shapley value has been commonly used in the field of the data economy as a quantitative and surrogate measure of data importance (i.e. Shapley value estimations can be used for data sampling, cleaning, pricing, abnormality detection etc): Naive computation of Data Shapley requires exponential time, and hence Monte Carlo (Ghorbani and Zou, 2019) and gradient-based methods are used to make it efficient (Jia et al., 2019). TracIn takes a similar idea of tracing the influence of individual training examples with gradient information. To make these algorithms practical and easy to use, DataScope (Karlaš et al., 2022) is developed as an end-to-end system that can efficiently compute the Shapley value of training data over the whole pipeline consisting of various ML algorithms and data transformation, making it a powerful tool for data debugging. With more mature data valuation, data providers are more motivated to contribute, fostering a more healthy and robust data-centric ecosystem.
@@
Model Combination Model combination (MC) strives to improve the overall system’s performance by either orchestrating or merging a series of (specialized) large models. The key benefits of model combination rely on the fact that there is usually little or no need for explicit training, and they can often result in better downstream performance and task-generalization capability. FrugalGPT (Chen et al., 2023i) routes quests in a cascading manner to different LLMs and uses a learned scoring function to decide whether to return the intermediate results in a flexible way, which drastically lowers the cost and improves the quality. Merging weights of multiple LLMs has been explored in many forms and shown to be effective. Popular methods include simple averaging (Wortsman et al., 2022), task arithmetic (Ilharco et al., 2023), multi-modal (encoders) merging (Wu et al., 2023b; Sung et al., 2023), merging based on learned routing function (Lu et al., 2023), SLERP, and weighted (conjugate gradient descent (Tam et al., 2023), stochastic and population-based optimization algorithm (Huang et al., 2024)) merging. MC can also be extremely promising for federated learning because only model weights are exchanged, and hence, data privacy is easier to guarantee. CoID Fusion (Don-Yehiya et al., 2023) proposes to collaboratively improve the multi-task learning of a base model by sending copies to workers and fusing the learned weights without data communication. The model combination can lead to compound systems [^4] https://bair.berkeley.edu/blog/2024/02/18/compound-ai-systems which consists of multiple LLMs working in synergy (via merging, routing, or knowledge sharing). AIOS (Mei et al., 2024) devises a mechanism to integrate multiple LLM agents into an operating system, the synergistic combination of which enables increasingly complex, multi-modal tasks that require reasoning, execution, and interaction with the physical world. Tandom Transformer (S et al., 2024) equips a smaller less accurate model with attention to the rich representation of a larger model that can process multiple tokens simultaneously, which serves as a stitched student-teacher system that improves both accuracy and efficiency in the downstream tasks. Developing complex compound systems poses several challenges: 1) how to co-optimize multiple LLMs, 2) identifying the failing (insecure) component is a lot harder than debugging a monolithic system, and 3) how to design mature data pipelines for different components of the large system. Addressing (some of) these problems will significantly increase the possibility of exciting new applications.
@@
Automation The increasing complexity of building a large foundation model requires a more mature automation process for democratization and agile development. AutoML (Hutter et al., 2019) has achieved remarkable success in many machine learning tasks over the last couple of years (Zimmer et al., 2020; Feurer et al., 2020; Erickson et al., 2020), which proves itself as a promising solution for large model automation. Applying AutoML techniques to LLMs, however, poses many challenges such as the cost for pretraining, the multitude of different stages, and performance indicators, making holistic optimization difficult or even infeasible (Tornede et al., 2024). Nonetheless, we will introduce some exemplary attempts targeting certain stages of the whole system. PriorBand (Mallik et al., 2023) tries to bridge the gap in the cost of Hyperparameter Optimization (HPO) between traditional ML and modern DL by utilizing expert beliefs and cheap proxy tasks. AdaBERT (Chen et al., 2021a) is an automated task-specific compression algorithm based on differentiable Neural Architecture Search (NAS) which is guided by both a task-oriented knowledge distillation loss and an efficiency-aware loss. To reduce the burden of prompt engineering, Automatic Prompt Engineer (APE) (Zhou et al., 2023c) proposes to leverage the interplay of several LLMs for automatic prompt generation and selection where one LLM proposes or modifies a prompt and another LLM rates it for selection. EcoOptiGen (Wang et al., 2023f) optimizes the utility and cost of decoding by finding better hyper-parameters, such as the number of responses, temperature, and max tokens, which demonstrates the potential of applying AutoML for the inference stage. One extremely exciting approach for building complicated and compound systems is to ask multiple LLMs to solve a big problem in a decomposed way cooperatively. One realization is to prompt the LLM or VLM to serve different purposes in a pipeline, which can be tremendously challenging to tune, optimize, modularize, and debug. DSPy (Khattab et al., 2023; 2022) tackles this by first separating the flow of the system from the parameters (i.e., model prompts and weights) at each step, and then dedicated algorithms are used to tune them with user’s defined metric. Even with all the works discussed above, the integration and development of Automated Large Models (AutoLM) has many challenges and opportunities simultaneously.